OpenStudy (anonymous):

write the equation of the line that passes through (3,8) and is parallel to the line -3x+y=12. Write your answer in slope intercept form

OpenStudy (anonymous):

Hi Welcome to Openstudy @mathsucks_0 :) First thing you need to know is the slope intercept form: \(\sf y=mx+b\) where m is the slop and b is the y-intercept. Next thing you need to do is determine the slope of -x+y=12, do you know how to do it?

OpenStudy (anonymous):

oops supposed to be ** -3x+y=12

OpenStudy (anonymous):

For slope I got -2/-3.. is that correct?

OpenStudy (anonymous):

Oh actually I was looking at the wrong problem! But I am not able to get the slope

OpenStudy (anonymous):

let's see.. i'll change it to slope intercept form, which will be: -3x+y=12 y=3x+12 m=3 and b=12 the intercept for this equation is 3. Now, since the line that you are looking for is PARALLEL to the given equation, they will have the SAME slope. So plug m=3 in the equation y=mx+b, what will you get for this part ?

OpenStudy (anonymous):

y=3x+b

OpenStudy (anonymous):

right :) Using your given point (3,8), we can now solve for the y-int or b in the equation. Plug in x=3 and y=8, then solve for b

OpenStudy (anonymous):

b=8/3?

OpenStudy (anonymous):

y=3x+b 8=3(3) + b -> plug in the points b= 8-3(3) -> evaluate to solve for b what will you get?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

I see what I did wrong, I missed where you wrote to plug in for x

OpenStudy (anonymous):

i'm glad that you are figuring out where you did wrong ^_^ that's a good habit for students! anyway, we're almost done so m=3, b= -1 y=mx+b, put the values of m and b in the equation and you'll get?

OpenStudy (anonymous):

y=3x-1

OpenStudy (anonymous):

and you're done ! :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome